Hydrogen peroxide (H2O2) is a powerful oxidizing agent used in concentrated solution in rocket fuels and in diluted solutions is hair bleach. An Aqueous solution of 25% (m/m) has a density of 1.05 g/ml. Calculate the concentration in terms of:



a. Mole Fraction of H2O2



b. Molarity


1
Expert's answer
2022-02-11T11:20:02-0500

Molality= m(solute)/m(solvent).p=m/V.Molality = 30/0.07= 428,571429.30/34 ...